The MacCauley Company has sales of $200 million and total operating expenses (excluding depreciation) of $130 million. Straight-line depreciation on the company's assets is $15 million over 4 years. The Upfront Cost equals to the total depreciable value on the company’s assets. Assume that all taxable income is taxed at 30 percent. Assume also that net operating working is 3% of sales in each year. Calculate the MacCauley Company's WACC using 4% cost of debt, 12% cost of equity. There is 45% equity and 55% debt are on the company’s balance sheet. What is NPV? Would you invest in the project? Is IRR greater or smaller than WACC? What is the payback period? (There is no salvage value at the end of 4 years)
Answer #1

Calculation of WACC

Weight of debt = 55%, Cost of debt = 4% , Weight of equity = 45% , Cost of Equity = 12%, Tax rate = 30%

WACC = Weight of debt x Cost of debt x (1- tax rate) + Weight of Equity x Cost of Equity = 55% x 4% x (1-30%) + 45% x 12% = 55% x 4% x 70% + 45% x 12% = 1.54% + 5.4% = 6.94%

Calculating upfront cost

Annual straight line depreciation = $15 million

Total upfront cost = No of years x Annual straight line depreciation = 4 x 15 = 60 million

Calculating investment in net working capital

Working capital required at beginning of year = Sales for the year x 3%

Working capital required in year 0 or beginning of year 1 = Sales for year 1 x 3% = 200 x 3% = $6 million

Incremental investment in working capital in year 0 =$6 million

Working capital required at beginning of year 2 or in year 1 = Sales in year 2 x 3% = 200 x 3% = $6 million

Incremental investment in working capital in a year = Working capital required in a year - Working capital required in previous year

Incremental investment in year 1 = Working capital required for year 1 - Working capital required for year 0 = 6 - 6 = 0 million

Negative value in year 4 represents recovery of cash from sale of working capital

Calculation of Incremental Investment in Working Capital
Year 0 1 2 3 4 5
Sales 0 200 200 200 200 0
Working Capital Required 6 6 6 6 0
Incremental Investment Working capital 6 0 0 0 -6

Calculating After tax operating cash flows

After tax operating cash flow = EBIT(1-tax rate) + Depreciation = (Sales - Cost - Depreciation)(1-tax rate) + Depreciation = (200 - 130 - 15)(1-30%) + 15 = 55(1-30%) + 15 = 38.50 + 15 = $53.50 million

Calculating Net cash flows

Calculating Net Cash Flows (in $ million)
Year 0 1 2 3 4
Initial Upfront Cost (a) -60
After tax operating Cash Flow (b) 53.50 53.50 53.50 53.50
Incremental Investment in Working Capital (c) 6 0 0 0 -6
Net Cash Flow = (a) + (b) - (c) -66 53.5 53.5 53.5 59.5
Year 0 1 2 3 4
Net Cash Flow (in $) -66000000 53500000 53500000 53500000 59500000

Calculating NPV of the project

NPV = Cash flow in year 0 + Sum of present values of cash flows for year 1 to 4 discounted at WACC

= -66000000 + 535000000 / ( 1 + 6.94%) + 535000000 / ( 1 + 6.94%)2 + 535000000 / ( 1 + 6.94%)3 + 595000000 / ( 1 + 6.94%)4

= -66000000 + 50028053.1139 + 46781422.3993 + 43745485.6923 + 45494222.4339 = 120049183.6394 = 120049183.64 (Rounded to two decimal places)

Since NPV of the project is greater than 0, therefore investment should be made in the project

Calculating IRR of the project

IRR of the project can be found out using IRR function in excel

Formula to be used in excel: =IRR(Cash flows)

60 Year 0 2 4 61 Net Cash Flow (in $) 6600000053500000535000005350000059500000, 62 IRR IRR(O61:S61 63

Using IRR function in excel, we get IRR of project = 72.689% = 72.69%

IRR is greater than WACC

Calculating Payback period

For calculating payback period, we will calculate cumulative cash flows

Cumulative cash flow in year 0 = Net cash flow in year 0 = -66000000

Cumulative cash flow in year = Cumulative cash flow for previous year + Net cash flow for a year

Cumulative cash flow for year 1 = cumulative cash flow for year 0 + net cash flow for year 1 = -66000000 + 53500000 = -12500000

Similarly cumulative cash flow can be found out for other years

Year 0 1 2 3 4
Net Cash Flow (in $) -66000000 53500000 53500000 53500000 59500000
Cumulative Cash Flows -66000000 -12500000 41000000 94500000 151000000

Cumulative cash flow change sign from negative to positive in year 2, therefore

Pay back period = 1 + (12500000 / 53500000) = 1 + (0.2336) = 1.2336= 1.23 years ( rounded to two decimal places)